girlcode:

An elephant can run one over four of a mile in 36 seconds. At this rate, which expression can be used to determine how fast an elephant runs in miles per hour? (5 points) a thirty-six seconds over one-fourth mile = 25 miles per hour b one-fourth mile over thirty-six seconds = 144 miles per hour c one-fourth mile over thirty-six seconds = 25 miles per hour d thirty-six seconds over one-fourth mile = 144 miles per hour

TWALKER16:

whitch do u think

girlcode:

I think it's c

jhonyy9:

ok so like a first step one hour has how many seconds ?

jhonyy9:

ok step by step one hour has how many minutes ? one minute has how many seconds ?

jakfishman:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 ok step by step one hour has how many minutes ? one minute has how many seconds ? \(\color{#0cbb34}{\text{End of Quote}}\) 1. 60 2. 60 is what you want right

girlcode:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jakfishman @girlcode u still need help??? \(\color{#0cbb34}{\text{End of Quote}}\) no sorry it took me so long to answer

jakfishman:

ok

jakfishman:

its fine

jhonyy9:

yes @jakfishman so @girlcode do you understand it and now than one hour has 60 minutes and one minute has 60 seconds so in this way how many seconds has one hour ?

jhonyy9:

courage please... this is easy

jhonyy9:

ok so use the rule of 3 simple and in this way you can write 1 minute has 60 sec 60 min. has x sec and now cross multiplie and calcule the x value x = 60*60/1 = 60*60 = ? sec

jhonyy9:

@girlcode ?

jhonyy9:

is offline

jhonyy9:

somebody knows the way how solve it ?

jhonyy9:

@supie i think you need to know - yes ?

jhonyy9:

courage i will help you - please collaborate